OpenStudy (anonymous):

What is 49a squared minus 169

OpenStudy (cwrw238):

use the identity a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

Can someone help me factor the quadratic equation 49a squared minus 168

OpenStudy (solomonzelman):

\[49a^2=(7a)^2\]\[169=13^2\]

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

My pleasure ! \(\Huge\color{blue}{ \sf ☼ }\)

OpenStudy (solomonzelman):

I can show you how to put up something on a keyboard if you want. A @swagapple123 do you want any symbols ?

OpenStudy (anonymous):

Ok sure

OpenStudy (anonymous):

Nvm I got it now but thx u are so nice

OpenStudy (solomonzelman):

1) Click and hold ALT 2) Click 7 5 3 It gives you: ± 1) Click and hold ALT 2) Click 7 5 9 It gives you: ± 1) Click and hold ALT 2) Click 2 or 1 It gives you: ☻ or ☺ (first is 2 and the second is 1) fr all those, use the numbers on the right of the keyboard, not the ones below F1, F2, F3...

OpenStudy (anonymous):

Thx

OpenStudy (solomonzelman):

You are welcome. I always liked symbols \(\Huge\color{lime}{♠ ♥ ♦ ♣} \)

OpenStudy (solomonzelman):

Good luck !

OpenStudy (anonymous):

☻ ☺ ♣ ◘ ○•♦♣♠☺☻♥

OpenStudy (solomonzelman):

\(\Huge\color{blue}{\bigstar~~ ▒ ~~~✂~~~▼~~~◙~~~◘~~~•} \)